Xylene, an organic compound that is a major component of many gasoline blends, contains carbon and hydrogen only. Complete combustion of a 17.12 mg sample of xylene in oxygen yielded 56.77 mg of CO2 and 14.53 mg of H2O. Determine the empirical and molecular formula of xylene. Molecular weight of xylene is 106.16 g/mol


1
Expert's answer
2021-09-21T01:55:37-0400

Xylene moles "=\frac{17.12}{106.16×1000}=0.00016moles"


Moles of "CO_2
=\frac{56.77}{44.01×1000}=0.0013"


Moles of "H_2O=
=\frac{14.53}{18.02×1000}=0.0008"


Moles ratios


"\frac{0.0013}{0.0008}=1.625"


"\frac{0.0008}{0.0008}=1"


Hence molecular fomula

The empirical formula is C 4H 5.

The molecular formula C8H10